OpenStudy (anonymous):

i am really having a har time dividing these polynomials 10a^3+19a+9/2a+3 can anyone help?

OpenStudy (anonymous):

This is a VERY useful site for learning this well. http://www.mathsisfun.com/algebra/polynomials-division-long.html

OpenStudy (amistre64):

you can split it up into component parts if you like :)\[\frac{10a^3}{2a+3} +\frac{19a}{2a+3}+\frac{3}{2a+3}\]

OpenStudy (amistre64):

the usual solution is long division...

OpenStudy (anonymous):

i tried doing long division, i keep getting crazy answers

OpenStudy (dumbcow):

\[5a ^{2}-7.5a+20.75-\frac{53.25}{2a+3}\]

OpenStudy (dumbcow):

did any of your crazy answers look like that

OpenStudy (anonymous):

no i didn't have decimals

OpenStudy (amistre64):

5a^2 -7.5a -------------------------- 2a+3 | 10a^3 +19a +9 -10a^3 -15a^2 ----------------- -15a^2 +19a +9 15a^2............ ya like the cows :)

OpenStudy (amistre64):

slow comp today.....:/